A rectangle is 3cm longer and 2cm less in width than a certain square. The area of the rectangle is 16cm^2 greater than the area of the square. What is the length of the rectangle?
Answer by JBarnum(2146)

(x+3)(x-2)=16+x^2
x^2+x-6=16+x^2
x-6=16
x=10
rectangle length=x+3
so 10+3=13